OpenStudy (anonymous):

Fan + Medal!!! Create your own polynomial with a degree of 3 that when divided by x + 2 has a remainder of -4

OpenStudy (cwrw238):

we can write it as (x -2)^3 - 4 = (x - 2)(x^2 -4x + 4) - 4 etc can you continue?

OpenStudy (cwrw238):

just multiply the second bracket by x then by -2 and simplify

OpenStudy (anonymous):

loll that should be x+2

OpenStudy (cwrw238):

oh heck yes sorry do the same with x + 2

OpenStudy (cwrw238):

(x + 2)(x^2 +4x + 4) - 4 = x^3 + 4x^2 + 4x + 2x^2 + 8x + 8 - 8 can you finish this off - just add like terms and you have it

OpenStudy (cwrw238):

thats -4 at the end - what the matter with me today!!

OpenStudy (anonymous):

so then just combine like terms and you get your polynomial?

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

okay :) Thanks so much!

OpenStudy (cwrw238):

you should get x^3 + 6x^2 + 12x + 4

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

so the answer would be x³ + 6x² + 12x - 4?

OpenStudy (anonymous):

@cwrw238

OpenStudy (cwrw238):

no +4 the end bit was +8 - 4

OpenStudy (cwrw238):

I made a mistake when i wrote + 8 - 8

OpenStudy (cwrw238):

ok? i must be tired i'm making silly mistakes all the time

OpenStudy (cwrw238):

x^3 + 6x^2 + 12x + 4 is correct

OpenStudy (anonymous):

oh okay! Thanks so much!!

OpenStudy (cwrw238):

yw